Charlotte earns time-and-a-half on

Saturdays and double time on

Sundays. She works 35 hours from

Monday to Friday, 8 hours on

Saturday and 5 hours on Sunday.

Calculate Charlotte’s total earnings if

her normal rate of pay is $15.40 per

hour.


1
Expert's answer
2021-03-16T07:16:15-0400

Normal pay (Monday to Friday) = 35 "\times" $15.40 = $539

Saturday pay = 8 "\times" $15.40 "\times" 1.5 = $184.8

Sunday pay = 5 "\times" $15.40 "\times" 2 = $154

Total pay = $877.80